A bag contains 52 tickets numbered 1,2...52 of which 5 are drawn at random and arranged in ascending order t1>t2>t3>t4>t5.find the probability that t4 being 42.
Answer by stanbon(48568)

A bag contains 52 tickets numbered 1,2...52 of which 5 are drawn at random and arranged in ascending order t1>t2>t3>t4>t5.find the probability that t4 being 42.
----------
To get the 42 on the 4th card you have to NOT get the 42 on the 1st, the
2nd, or the 3rd card. Once you get the 42 on the 4th card the next card
can be any of the remaining cards.
Prob(no 42 on 1st draw)= 51/52
Prob(no 42 on 2nd draw) = 50/51
Prob(no 42 on 3rd draw) = 49/50
Prob(42 on the 4th draw) = 1/49
Prob(any card on the 5th draw) = 48/48
-----------------
P(42 on 4th draw) = (51/52)(50/51)(49/50)(1/49)(48/48)
= (1/52)(1/49)
= 1/2548
==========
Cheers,
Stan H.